How many pounds of pure peanuts must be combined with a 20% peanut mix to produce 2 pounds of a 50% peanut mix?
Found 2 solutions by addingup, MathTherapy:
Answer by addingup(3677)

Let pure peanuts be x. Then, 20% peanuts will be 0.2x
0.2x + 1(2-x) = 0.50(2)
0.2x + 2-x = 1
-0.8x = -1
x = 1.25

Answer by MathTherapy(10552)


How many pounds of pure peanuts must be combined with a 20% peanut mix to produce 2 pounds of a 50% peanut mix?
Pure peanut is 100%
Let amount of pure peanut be P
Then amount of 20% peanut = 2 -P
We then get: 1P + .2(2 - P) = .5(2)
P + .4 - .2P = 1
.8P = 1 - .4
Amount of pure (100%) peanut, or
As the resultant (50%) is FARTHER from pure (100%) peanut than it is to 20% peanut mix, it's expected that
the amount of 100% (pure) peanuts will be less than the amount of 20% peanut.